# Minimax Decision Trees via Martingale Approximations

### v1

@@ Abstract
We develop a martingale-based approach to constructing decision trees that efficiently approximate a target variable through recursive conditioning. We introduce MinimaxSplit, a novel splitting criterion that minimizes the worst-case variance at each step, and analyze its cyclic variant, proving an exponential error decay rate under mild conditions. Our analysis builds upon partition-based martingale approximations, providing new insights into their convergence behavior. Unlike traditional variance-based methods, MinimaxSplit avoids end-cut preference and performs well in noisy settings. We derive empirical risk bounds and also explore its integration into random forests.
Keywords: martingale approximation; convergence rates; tree-based models; non-parametric regression

### v2

# Stabilizing the Splits through Minimax Decision Trees

@@ Abstract
By revisiting the end-cut preference (ECP) phenomenon associated with a single CART (Breiman et al. (1984)), we introduce MinimaxSplit decision trees, a robust alternative to CART that selects splits by minimizing the worst-case child risk rather than the average risk. For regression, we minimize the maximum within-child squared error; for classification, we minimize the maximum child entropy, yielding a C4.5-compatible criterion. We also study a cyclic variant that deterministically cycles coordinates, leading to our main method of cyclic MinimaxSplit decision trees.
We prove oracle inequalities that cover both regression and classification, under mild marginal non-atomicity conditions. The bounds control the tree’s global excess risk by local worst-case impurities and yield fast convergence rates compared to CART. We extend the analysis to ensembles that subsample coordinates per node.
Empirically, (cyclic) MinimaxSplit trees and their forests improve on structured heterogeneity data such as EEG amplitude regression over fixed time horizons, seasonal air quality forecasting, and image denoising framed as non-parametric regression on spatial coordinates.
Keywords: CART; convergence rates; decision tree models; non-parametric regression